# with Controllable Camera and Object Motions

### v2

@@ 4 The Proposed Approach
Given $N$-length camera poses $\mathcal{C}_{RT}={RT}_{cam}^{1:N}$, object poses $\mathcal{O}_{RT}=\left\{RT_{obj_{i}}^{1:N}\right\}_{i=1}^{N_{o}}$ of $N_{o}$ objects in the global coordinate system, and content description $\mathbf{C}_{p}$, we aim to generate the video that reveals correct motion in real world.
As shown in Table 2, most methods cannot independently or jointly control object and camera movements. For example, AnimateDiff [9] and CameraCtrl [10] only support camera control, while other methods [43, 54, 50, 46] handle image-space trajectories for object motion control, often facing motion entanglement issues. Although Direct-a-Video [50] introduces several camera types and allows explicit control, it is limited to simple motion patterns. MotionCtrl [46], the closest to ours, trains two motion modules separately but lacks comprehensive pose annotations, resulting in suboptimal simultaneous control of camera and object motions. MotionCtrl’s object motion control module can only handle image-space trajectory, without accounting for orientation and distance. Additionally, MotionCtrl simply applies standard diffusion loss [11] when training motion modules, which further hinders its ability to disentangle camera and object motions within a video. FMC trained on SynFMC introduces Camera Motion Controller (CMC) and Object Motion Controller (OMC) to address these limitations. Among them, OMC receives 6D pose and coarse mask of the object to perceive its spatial location and orientation, thereby achieving a realistic appearance from various viewpoints. Additionally, the training objectives enable FMC to disentangle the motion effects of objects and the camera in the video, allowing independent or simultaneous control of camera and object motions.
$\bullet$ Preliminary. 1) During training, T2V diffusion models [9, 12, 59, 56, 36, 13] add Gaussian noise $\epsilon$ to image sequence $\mathbf{z}_{0}^{1:N}$, resulting in noisy latents $\mathbf{z}_{t}^{1:N}$ at $t$ time step. Network $\varepsilon_{\theta}$ then is trained to infer the injected noise from current latents. 2) AnimateDiff [9] uses Domain Adapter to adapt spatial modules to low-quality content. We apply this to bridge the synthetic-real video domain gap. 3) Following CameraCtrl [10], we use plücker embedding to represent camera pose for geometric interpretation.
@@ 4.1 Free-Form Motion Control
The overall architecture is shown in Figure 5. We train the proposed FMC (Free-Form Motion Control) in 3 stages. First, LoRA modules [15] are injected into the spatial blocks as Domain Adapter [9], learning the rendered content. The temporal modules remain inactive at this stage, and images are randomly sampled from synthetic data. Next, CMC parameters are updated to learn camera motion, with LoRA weights from the previous step loaded and temporal modules introduced to the model. Finally, OMC is trained to decouple object dynamics from camera motion, while other parameters are frozen.
$\bullet$ Construction of Training Data. We randomly sample frames from the original video at uniform intervals and obtain the corresponding object masks. Since SynFMC has detailed descriptions of the background, objects and their actions for each time step, video text descriptions can be automatically generated by specific templates. Then, these labels with pose information of camera and objects are used for training motion controllers.
$\bullet$ Camera Motion Controller (CMC). It consists of two parts as shown in Figure 5. Camera Encoder receives the initial camera pose (translation values are set to 0) and plücker embeddings of subsequent frames. The introduction of the initial pose helps to determine the perspective at the start time. Then, the outputs are processed by Camera Adapter to modulate the features in temporal blocks. Due to the dynamic of the background being only affected by camera motion, camera loss $L_{cam}$ is applied in this stage:
@@
Figure 5: The architecture of FMC. In the first stage, we randomly sample the images from synthetic videos and update the parameters from injected Domain Adapter. Next, the modules from CMC are learned. It consists of two parts: Camera Encoder and Camera Adapter, where the Camera Adapter is introduced into the temporal modules. Finally, we train the Object Encoder from OMC. It receives the object pose features, which are repeated in the corresponding object region. We use Gaussian blur kernel centered at the centroid to prevent the need of precise masks. Then, the output is multiplied by the coarse masks to modulate the features in the main branch.
@@ 5 Experiments
$\bullet$ Implementation Details. The proposed FMC is based on AnimateDiff V3 [9], trained with 16-length $256\times 384$ videos, Adam optimizer with a learning rate of $1e^{-4}$. Domain Adapter is trained with 8K iterations in a batch size of 128. CMC and OMC are trained with 50K iterations with a batch size of 8. $\lambda_{c}$ and $\lambda_{o}$ are set to 0.6 and 0.3.
@@
Figure 6: Independent control over camera and object motions. The results in (a) reveal that all methods [10, 46] effectively reflect the input conditions. For object motion control, the compared methods [46, 50] fail to maintain a stationary camera. As shown in green boxes from (b), the camera effect causes the dynamic in the background (e.g., movement of flower in row 5).
@@ 5.1 Comparisons with State-of-the-Art Methods
We first compare independent control over camera motion and object motion, as in previous methods [46, 10]. Then, we demonstrate FMC ’s superior performance in simultaneous control. Finally, we showcase additional examples across different scenes to validate the effectiveness of SynFMC and FMC.
$\bullet$ Independent Control of Camera Motion. MotionCtrl [46] and CameraCtrl [10] are selected for this comparison as they accept explicit camera information. In Figure 6(a), we simulate two camera motions and scale the translation to fit the input range required by these methods. FMC and the compared methods effectively reflect the input conditions. The CamTransErr & CamRotErr in Table 3 also show that FMC achieves comparable results in controlling camera.
@@
$\bullet$ Independent Control of Object Motion. For object control, we compare with MotionCtrl [46] and Direct-a-Video [50]. For these methods, we project the global trajectory into image space using camera and object pose information. For our FMC, the camera pose condition is set to identity matrices as static camera. As shown in Figure 6(b), the compared methods fail to maintain a stationary camera (showing dynamic movement in the background), indicating that image space trajectories entangle the object and camera motions. For example, the 2nd example of Figure 6(b), the result of Direct-a-Video [50] shows the change of flower location, which is caused by dynamic camera. In contrast, our method effectively alleviates this issue by incorporating static camera poses as constraints.
$\bullet$ Simultaneous Control of Camera and Object Motions. We explore the combination of both camera and object control signals. Since Direct-a-Video [50] supports only basic camera motion, we choose MotionCtrl [46] as the comparison method. We randomly simulate movements for both the object and the camera, resulting in varied image-space trajectories. As shown in Figure 7, videos generated by FMC more faithfully align with the specified conditions. While MotionCtrl captures the camera’s motion, it struggles to generate realistic object dynamics. These results demonstrate the effectiveness of our method in achieving simultaneous control of camera and object motions. The object error metrics in Table 3 show that our method achieves better results in object motion control. Additionally, FMC achieves higher scores in user study, as shown in Table 4, outperforming previous methods in quality and fidelity for both camera and object motions.
Figure 8 shows video generation results across 4 different cases: static single-object, dynamic single object, static multi-object, and dynamic multi-object. Thanks to the diversity and complexity of camera and motion motion patterns in SynFMC, FMC effectively learns a range of diverse, advanced, and complex shots. In the 2nd row of Figure 8, for example, the camera initially captures the person from the front and then follows from behind. The last two rows demonstrate the performance in multi-object scenarios, where the relative motion between objects and the camera conforms closely to the input conditions.
Table 4: User study in quality, text similarity, and motion accuracy.
@@ 5.2 Ablation Study
$\bullet$ SynFMC Dataset. To validate the effectiveness of complete camera and object pose annotations in SynFMC and evaluate the dataset generalization, we train MotionCtrl [46] on it, adapting the annotations to its input format. We first train the camera module, then optimize the object module in two ways: without camera poses, as in [46], and with camera poses, as our FMC. As shown in Figure 9, incorporating known camera poses when optimizing object module allows the object to follow the input trajectory more accurately, reducing the risk of it leaving the field of view. Object motion errors in Table 5 further highlight the benefits of using complete camera and object pose annotations.
$\bullet$ OMC. As shown in the first two rows of Figure 10 and object motion errors in Table 5, FMC outperforms MotionCtrl [46] trained on SynFMC in motion accuracy. This improvement is brought by OMC’s ability to process 6D poses, generating more realistic object appearances based on orientation and object’s distance from the camera, reflected by the size of coarse masks. In contrast, MotionCtrl’s object motion control module can only handle 2D image-space trajectories without pose and distance information, limiting alignment accuracy with the input.
@@
$\bullet$ Training Objectives. We conduct two experiments to assess the impact of $L_{cam}$ in Eq. 1 and $L_{obj}$ in LABEL:eq:obj-ddpm-loss. First, we train CMC using only the standard diffusion loss [11, 44]. As shown in the 3rd row of Figure 10, model without $L_{cam}$ tends to shift foreground objects to achieve similar relative motion, which does not accurately match the input pose. The camera motion error in Table 5 underscores the effectiveness of $L_{cam}$. Similarly, training OMC without $L_{obj}$ leads to undesired object appearances, as shown in the 5th row of Figure 10, with object motion error in Table 5 confirming the benefit of $L_{obj}$.
$\bullet$ Adaptability with Different T2I Personalized models. As shown in Figure 2, FMC can be adapted to various personalized backbones [38, 41, 29], showing that our proposed dataset SynFMC and corresponding training strategy do not impair the model’s original generative capabilities.
@@ 6 Conclusion
This work introduces SynFMC, a dataset with comprehensive pose information and diverse assets, offering both standard and complex shots that are difficult to capture in real life, with trajectories resembling real-world scenarios. With SynFMC, the proposed method FMC enables independent or joint control of object and camera motions within a single video. Experimental results demonstrate the effectiveness of both SynFMC dataset and FMC.

### v3

@@ 4 The Proposed Approach
Given $N$-length camera poses $\mathcal{C}_{RT}$ = ${RT}_{cam}^{1:N}$, object poses $\mathcal{O}_{RT}$ = $\{RT_{obj_{i}}^{1:N}\}_{i=1}^{N_{o}}$ of $N_{o}$ objects in the global coordinate system, and content description $\mathbf{C}_{p}$, we aim to generate the video that reveals correct motion in real world.
As shown in Tab. 2, most methods cannot independently or jointly control object and camera movements in a 3D-aware manner. For example, AnimateDiff [11] and CameraCtrl [12] only support camera control. Methods [44, 54, 51, 47] using image-space trajectories faces motion entanglement issues and can’t control the orientation. Although Direct-a-Video [51] introduces several camera types and allows explicit control, it is limited to simple motion patterns. MotionCtrl [47], the closest to ours, trains two motion modules separately but lacks comprehensive 6D pose annotations, resulting in suboptimal simultaneous control of camera and object motions. Additionally, it simply applies standard diffusion loss [13] when training motion modules, which further hinders its ability to disentangle camera and object motions within a video. FMC trained on SynFMC introduces Camera Motion Controller (CMC) and Object Motion Controller (OMC) to address these limitations. OMC receives 6D pose and coarse mask of the object to perceive its spatial location and orientation, achieving a realistic appearance from various viewpoints. The training objectives enable FMC to disentangle the motion effects of objects and the camera in the video, allowing independent or joint 3D-aware control of camera and object motions.
$\bullet$ Preliminary. 1) T2V diffusion models [11, 14, 58, 56, 36, 15] add Gaussian noise $\epsilon$ to image sequence $\mathbf{z}_{0}^{1:N}$ in training, resulting in noisy latents $\mathbf{z}_{t}^{1:N}$ at $t$ time step. Network $\varepsilon_{\theta}$ then is trained to infer the injected noise from current latents. 2) LoRAs [17, 11] are used to learn different content styles. We apply this to bridge the synthetic-real video domain gap. 3) Following [12], we use plücker embedding to represent camera pose for geometric interpretation.
Figure 6: Domain LoRA. We sample the first frame of generated videos under without and with Domain LoRA settings.
Figure 7: The architecture of FMC. In the first stage, we randomly sample the images from synthetic videos and update the parameters from injected Domain LoRA. Next, the modules from CMC are learned. It consists of two parts: Camera Encoder and Camera Adapter, where the Camera Adapter is introduced into the temporal modules. Finally, we train the Object Encoder from OMC. It receives the 6D object pose features, which are repeated in the corresponding object region. We use Gaussian blur kernel centered at the centroid to prevent the need of precise masks. Then, the output is multiplied by the coarse masks to modulate the features in the main branch.
@@ 4.1 Free-Form Motion Control
Fig. 7 shows the overall architecture of the proposed FMC method. We train it in 3 stages. First, Domain LoRA [17] are injected into spatial blocks to adapt to rendered content, with temporal modules inactive and images randomly sampled from synthetic data. Fig. 6 shows the effectiveness of this stage in bridging the domain gap. Then, CMC is trained to learn camera motion, introducing temporal modules and loading LoRA from the previous step. Finally, OMC is trained to decouple object dynamics from camera motion, with other parameters frozen. During inference, LoRA modules are dropped to maintain the quality of base model.
$\bullet$ Camera Motion Controller (CMC). It consists of two parts as shown in Fig. 7. Camera Encoder receives plücker embeddings, where the initial camera pose (translation values are set to 0) and the relative camera poses are used for the first and subsequent frames respectively. The initial pose helps to determine the perspective at the start time. Then, the outputs are processed by Camera Adapter to modulate the features in temporal blocks. Due to the dynamic of the background being only affected by camera motion, camera loss $L_{cam}$ is applied in this stage:
@@ 5 Experiments
$\bullet$ Implementation Details. The proposed FMC is based on AnimateDiff V3 [11], trained with 16-length $256\times 384$ videos, Adam optimizer with a learning rate of $1e^{-4}$. Domain Adapter is trained with 8K iterations in a batch size of 128. CMC and OMC are trained with 50K iterations with a batch size of 8. $\lambda_{c}$ and $\lambda_{o}$ are set to 0.6 and 0.3. In experiments, we find that 26K data samples suffice for our method to learn 3D-aware motion control.
@@
Figure 8: Independent controls over camera and object motions. Results in (a) reveal that all methods [12, 47] effectively reflect the camera conditions. For object motion, the compared methods [47, 51] fail to maintain a stationary camera as shown in green boxes from (b) (e.g., movement of flower in row 5). Furthermore, they also present low fidelity of object orientation (3D axes in conditions).
@@ 5.1 Comparisons with State-of-the-Art Methods
We first compare independent controls over camera motion and object motion with previous methods [47, 12]. Then, we demonstrate FMC ’s superior performance in simultaneous control. Finally, we showcase additional examples across different scenes to validate the effectiveness of SynFMC and FMC. For fairness, we compare with U-Net based methods. More examples are in supplementary.
$\bullet$ Independent Control of Camera Motion. MotionCtrl [47] and CameraCtrl [12] are selected for this comparison as they accept explicit camera information. In Fig. 8(a), we simulate two camera motions and scale the translation to fit the input range required by these methods. FMC and the compared methods effectively reflect the input conditions. The CamTransErr & CamRotErr in Tab. 3 also show that FMC achieves comparable results in controlling camera.
@@
$\bullet$ Independent Control of Object Motion. For object control, we compare with MotionCtrl [47] and Direct-a-Video [51]. For these methods, we project the global trajectory into image space using camera and object pose information. As shown in Fig. 8(b), the compared methods fail to maintain a stationary camera (showing dynamic movement in the background), indicating that image space trajectories entangle the object and camera motions. For example, the 2nd example of Fig. 8(b), Direct-a-Video [51] shows the change of flower location caused by dynamic camera. Our method effectively alleviates this issue by incorporating static camera poses as constraints. Furthermore, since OMC receives 6D pose of objects, FMC can achieve high fidelity of object orientation with input condition.
Table 4: User study in quality, text similarity, and motion fidelity.
@@
$\bullet$ Simultaneous Control of Camera and Object Motions. We explore the combination of both camera and object control signals. Since Direct-a-Video [51] supports only basic camera motion, we choose MotionCtrl [47] as the comparison method. We randomly simulate movements for both the object and the camera, resulting in varied trajectories. As shown in Fig. 9, videos generated by FMC more faithfully align with the specified conditions. While MotionCtrl captures the camera’s motion, it struggles to generate realistic object dynamics. These results demonstrate the effectiveness of FMC in achieving simultaneous control of camera and object motions. The object error metrics in Tab. 3 show that our method achieves better results in object motion control. Additionally, FMC achieves higher scores in user study as shown in Tab. 4, outperforming previous methods in quality and motion fidelity.
Fig. 10 shows video generation results across 4 different cases: static single-object, dynamic single-object, static multi-object, and dynamic multi-object. Thanks to the diversity of motion patterns in SynFMC, FMC effectively learns a range of diverse, advanced, and complex shots. In the 2nd row of Fig. 10, for example, the camera initially captures the person from the front and then follows from behind. The last two rows demonstrate the performance in multi-object scenarios, where the relative motion between objects and the camera conforms closely to the input conditions.
5.2 Ablation Study
@@
$\bullet$ SynFMC Dataset. To validate the effectiveness of complete camera and object pose annotations in SynFMC and evaluate the dataset generalization, we train MotionCtrl [47] on it, adapting the annotations to its input format. We first train the camera module, then optimize the object module in two ways: without camera poses, as in [47], and with camera poses, as our FMC. As shown in Fig. 11, incorporating known camera poses when optimizing object module allows the object to follow the input trajectory more accurately, reducing the risk of it leaving the field of view. Object motion errors in Tab. 5 further highlight the benefits of using complete camera and object pose annotations.
$\bullet$ OMC. As shown in the first 2 rows of Fig. 12 and object motion errors in Tab. 5, FMC outperforms MotionCtrl [47] trained on SynFMC in motion accuracy. This improvement is brought by OMC’s ability to process 6D poses, generating more realistic object appearances based on orientation and object’s distance from the camera, reflected by the size of coarse masks. [47]’s object motion control module can only handle 2D image-space trajectories without pose and distance information, limiting alignment accuracy with input.
@@
$\bullet$ Training Objectives. We conduct two experiments to assess the impact of $L_{cam}$ in Eq. 1 and $L_{obj}$ in LABEL:eq:obj-ddpm-loss. First, we train CMC using only the standard diffusion loss [13, 45]. As shown in the 3rd row of Fig. 12, model without $L_{cam}$ tends to shift foreground objects to achieve similar relative motion, which does not accurately match the input pose. The camera motion error in Tab. 5 underscores the effectiveness of $L_{cam}$. Besides, training OMC without $L_{obj}$ leads to undesired object appearances, as shown in the 5th row of Fig. 12, with object motion error in Tab. 5 confirming the benefit of $L_{obj}$.
$\bullet$ Adaptability with Different T2I Personalized models. As shown in Fig. 2, FMC can be adapted to various personalized backbones [38, 41, 30], showing that our proposed dataset SynFMC and corresponding training strategy do not impair the model’s original generative capabilities.
@@ 6 Conclusion
This work introduces SynFMC, a dataset with comprehensive 6D pose information and diverse assets, offering both standard and complex shots that are difficult to capture in real life, with trajectories resembling real-world scenarios. With SynFMC, the proposed method FMC enables independent or joint 3D-aware control of object and camera motions within a single video. Experimental results demonstrate the effectiveness of both SynFMC dataset and FMC.
@@
Acknowledgement.
This project was supported by the National Natural Science Foundation of China (NSFC) under Grant No. 62472104. This work was supported by Damo Academy through Damo Academy Innovative Research Program. Dr Tao’s research is partially supported by NTU RSR and Start Up Grants. This work was in part supported by National Natural Science Foundation of China (Grant No. 62276067).